OpenStudy (anonymous):

If (x – 5) is a factor of f(x), which of the following must be true?

OpenStudy (anonymous):

A root of f(x) is x = –5. A root of f(x) is x = 5. Both x = –5 and x = 5 are roots of f(x). Neither x = –5 nor x = 5 is a root of f(x).

OpenStudy (anonymous):

lease help p

OpenStudy (anonymous):

please

OpenStudy (zarkon):

if \((x-5)\) is a factor of \(f(x)\) then \(f(x)=(x-5)g(x)\) for some function \(g(x)\) provided that \(g(x)\) is defined at 5 you get \(f(5)=(5-5)g(5)=0g(5)=0\)

OpenStudy (anonymous):

ty c right i think i miscalculated